OpenStudy (campbell_st):

Pythagoras arranges the numbers 1, 2, 3, ..., 11, 12 into 6 pairs so that the sum of any 2 numbers in a pair is prime and no two of these primes are equal. Find the primes and the pairs

OpenStudy (zarkon):

5,7,11,13,19,23

OpenStudy (zarkon):

(1,4)(2,5)(3,8)(6,7)(9,10)(11,12)

OpenStudy (kainui):

I'm curious how people are getting to the answer, this is interesting. Since all pairs are greater than 2, then all the pairs should be odd. So that means each pair is an even with an odd. But after that I am not sure where to go.

OpenStudy (anonymous):

(1,2)=3 (3,4)=7 (5,6)=11 (7,8)=13 (10,9)=19 (11,12)=23

OpenStudy (zarkon):

look at the sum of 1 through 12

ganeshie8 (ganeshie8):

I'd start by dividing them into two groups : odd, even you need to pick numbers from different groups 1 2 3 4 5 6 7 8 9 10 11 12

OpenStudy (campbell_st):

is it the only solution...

OpenStudy (anonymous):

seems like not unique

OpenStudy (campbell_st):

@ marki, you have an error 7 + 8 = 15 you listed it as 13

ganeshie8 (ganeshie8):

trial and error is the only way i guess..

OpenStudy (mathstudent55):

@Marki: 7 + 8 = 15 (not prime)

OpenStudy (anonymous):

hehe

OpenStudy (anonymous):

sorry

OpenStudy (kainui):

It appears as though there are 6! possiblities... Who would like to sort through this?

ganeshie8 (ganeshie8):

6*6 = 36 cases to check

OpenStudy (campbell_st):

nice job @Zarkon , nothing like the old guess and check method

OpenStudy (anonymous):

aha !

OpenStudy (campbell_st):

nope only 2 solutions... I've found...

ganeshie8 (ganeshie8):

I think there is no standard method(formula) to find when a sum of two integers equals prime

OpenStudy (kainui):

There are only 8 possible primes we can make, maybe we can look at it going backwards by looking at constructing from the range of 3 to 23?

OpenStudy (zarkon):

(2,3)(1,6)(4,7)(5,8)(9,10)(11,12)

OpenStudy (campbell_st):

that's it @Zarkon has both solutions...

OpenStudy (campbell_st):

hope it made you think..

OpenStudy (zarkon):

and it only took me 3 guesses ;)

OpenStudy (zarkon):

after narrowing down the possibilities

OpenStudy (kainui):

The sum of these pairs is the sum of all the numbers which is the sum of primes, so we have 12(13)/2 =6*13=78 so we can find whatever sums of the 8 primes is 78 and use them.

ganeshie8 (ganeshie8):

Clever !

OpenStudy (kainui):

So we know: 98-p-q=78 so we have: p+q=20 Find the two primes that sum to 20 and you just use the other 8 ;)

OpenStudy (zarkon):

that is one of the things I did "look at the sum of 1 through 12"

OpenStudy (kainui):

So you can exclude 17 and 3 for one possible way or exclude 13 and 7 for the other way, and no others exist.

ganeshie8 (ganeshie8):

this question reminds me of goldbach's conjecture

OpenStudy (anonymous):

1 2 3 4 5 6 7 8 9 10 11 12 (1,2) (1,4) ( 1,6) (1,8) (1,10) (1,12) ---- > 3,5,7,9(ignore ) ,11,13 (3,2) (3,4) (3,6) (3,8) (3,10) (3,12)----> 5,7,9(ignore),13,15(ignore (5,2) (5,4) (5,6) (5,8) (5,10) (5,12)---->7,9(ignore) ,11,13,15(ignore),17 (7,2) (7,4) (7,6) (7,8) (7,10) (7,12)---->9(ignore) ,11,13,15(ignore) ,17,19 (9,2) (9,4) (9,6) (9,8) (9,10) (9,12)---->11,13,15(ignore) ,17,19,23 (11,2) (11,4) (11,6) (11,8) (11,10) (11,12) ---> 13,15,17,19,21(ignore),23

OpenStudy (anonymous):

lol long time for posting :(

ganeshie8 (ganeshie8):

goldbach conjecture is about representing integers as sum of two primes i guess

ganeshie8 (ganeshie8):

lol Zarkon said he was able to figure it out by just 3 guesses im still wondering how to reduce 36 to 3

OpenStudy (anonymous):

aha yeah , i tout it to my friend last day and found that this is lema the origin is like this :-

OpenStudy (anonymous):

i'll tell u how ,least integer is 3 , highest is 23 so we have 3,5,7,11,13,17,19,23 btw ( 8 primes ) hmm then wake take reference lol not sure if he did the same

OpenStudy (anonymous):

hmm :|

OpenStudy (kainui):

Yeah that will work @Marki that's what I did. It turns out that there are two possible pairs of primes you can exclude, but once you start to fill out a chart by process of elimination starting at the smallest, if you exclude 7 and 13 from the list of 8 you have to make 3 with (1,2) but then you can't make 5 because (3,4) is the smallest pair you have left.

OpenStudy (campbell_st):

hope it made you think

OpenStudy (kainui):

Yeah thanks I enjoyed it!